one month keisha rented 5 movies and 3 video games for a total of $41. The next month she rented 7 movies and 9 videos games for a total of $85. FInd the rental cost for each movie and each video game
Answer by fractalier(6550)

Call their prices m and v. We have
5m + 3v = 41
7m + 9v = 85
Now multiply the top one by three and subtract it from the bottom one...
7m + 9v = 85
-(15m + 9v = 123)
-----------------
-8m = -38
m = 38/8 = $4.75 so that
5(4.75) + 3v = 41
23.75 + 3v = 41
3v = 17.25
v = $5.75